OpenStudy (anonymous):

(square root 48 + square root 27)2

OpenStudy (aum):

\((a+b)^2 = a^2 + 2ab + b^2\)

OpenStudy (anonymous):

\[\sqrt{48}\]= 6.92..... + \[\sqrt{27}\]= 5.196.....= 12.124... * 2 = 24.25.

OpenStudy (aum):

\[ (\sqrt{48} + \sqrt{27})^2 = (\sqrt{48})^2 + 2(\sqrt{48})(\sqrt{27}) + (\sqrt{27})^2 = \\ 48 + 2\sqrt{48*27} + 27 = 75 + 2\sqrt{16*3*3*9} = \\ 75 + 2\sqrt{4^2*3^2*3^2} = 75 + 2 * 4 * 3 * 3 = 75 + 72 = 147\\ \]

jhonyy9 (jhonyy9):

sqet48 = 4sqrt3 sqer27 = 3sqrt3 assum it will get 7sqrt3 so 7sqrt3 squared is equal 49*3 = 147 hope this will help you